# Coherent Local Explanations for Mathematical Optimization

### v1

@@ 4.4 Vehicle Routing Problem
We consider the Capacitated Vehicle Routing problem (CVRP). An instance of this problem is given by a complete graph $G=(V,A)$ where $V$ consists of a depot node $v_{0}$ and $n$ client nodes each with a corresponding demand $d_{j}$. Moreover, each arc $(v_{j},v_{k})$ has associated costs $c_{jk}$. Lastly, there are $m$ vehicles, each with a capacity of $M$. The goal of the problem is to find at most $m$ routes of minimum costs such that each route starts and ends at the depot, each client is visited exactly once, and the total demand on each route does not exceed the vehicle capacity. To formulate $R_{C}$, we use the Miller-Tucker-Zemlin formulation for the CVRP which can be found in Equation 14 in the appendix. The decision variables are then denoted by $x_{jk}$ and equal $1$ if arc $(v_{j},v_{k})$ is used in the solution and 00 otherwise.
For this experiment, we consider the parameter vector consisting of the demands $\bm{d}$ and the costs of the arcs from the clients to the depot $\bm{c_{0}}$, i.e., $\bm{\theta}=(\bm{d},\bm{c_{0}})$. The present problem has symmetric costs and consists of 16 clients and 4 vehicles. As a solver for this NP-hard problem, we let the Google OR-Tools heuristic search for a solution for 5 seconds (Furnon & Perron, ). We aim for explanations for the objective value, and which clients are visited before returning to the depot, i.e., the decision variables $x_{j0}$ for $j\in[n]_{1}$.
As shown in Table 4 the explanation found by CLEMO is significantly more coherent without considerably losing in accuracy compared to the LR benchmark. We visualize the explanations found by CLEMO in Figure 3. In this figure, we first see the solution to the present problem found by Google OR-Tools in gray. Next, the feature contribution of the demands and costs are visualized via node and edge colors respectively. Combined with an overview of the top 10 most contributing features, this shows which features are the key components influencing the objective value as found by the solver. Thus, Figure 3 tells us that the objective value is mainly affected by the distance towards the nodes far from the depot. In Figure 9 in the appendix, we present an additional explanation for the decision variable $x_{20}$.
@@
value
Decision
vector
Objective
Feasible
region
@@
5 Conclusion & Further Research
In this paper, we propose CLEMO, a sampling-based method that can be used to explain arbitrary exact or heuristic solution algorithms for optimization problems. Our method provides local explanations for the objective value and decision variables of mathematical optimization models. Contrary to existing methods, CLEMO enforces explanations that are coherent with the underlying model structure which enhances transparent decision-making. By applying CLEMO to various optimization problems we have shown that we can find explanations that are significantly more coherent than benchmark explanations generated using LIME without substantially compromising prediction accuracy.
This work focuses on explaining the objective value and decision variables. However, one could easily extend the concept to explanations of other components such as constraint slacks, runtime, optimality gap, etc. Another extension of our work could be to consider other types of interpretable functions such as decision trees. Lastly, CLEMO could be a useful method to explain synergies between ML and optimization models, e.g., in predict-then-optimize models.
@@ A.1 Mathematical Models of The Optimization Problems
In this section, we present the formulation of the Shortest Path problem (13) and the Capacitated Vehicle Routing problem (14). For the latter, we use the Miller-Tucker-Zemlin formulation as described by Kara et al. (2004).
@@ A.3 Coherence for Linear Optimization Problems
$\min_{\bm{\beta}_{x_{j}}}\ \sum_{i=0}^{N}w^{i}\|\bm{x}^{i}_{j}-\bm{\beta}_{x_{j}}^{\intercal}\bm{\theta}^{i}\|^{2}\quad j=1,\ldots,p.$
fulfill the coherence Condition (3).
@@
$\bm{\beta}_{f}^{*}=(\bm{\Theta}^{\intercal}\bm{W}\bm{\Theta})^{-1}\bm{\Theta}^{\intercal}\bm{W}\bm{y^{f}}$
@@
$=\sum_{j=1}^{p}\hat{c}_{j}\bm{\theta}^{\intercal}\bm{\beta}_{x_{j}}$
@@ A.4.2 Knapsack Problem
Next to the standard deviation of feature contribution, we consider an additional measure for stability, the feature stability index (FSI). This is an adaptation of the variables stability index (VSI) as presented in (Visani et al., 2022). The higher this measure, the more the non-zero features found by the different models due to resampling overlap. For a consistent explanation, the overlap should be large. As we apply CLEMO on 10 different datasets for each instance of each type of knapsack problem, we obtain 10 surrogate models given by $\bm{\beta}^{1}_{CL},\dots,\bm{\beta}^{10}_{CL}$. We denote $\mathcal{F}_{k,j}^{i}$ for the set of the top-$k$ most contributing, non-zero features of the $j$-th component of $\bm{\beta}^{i}_{CL}$. We define the $(k,j)$-concordance of two models $\bm{\beta}^{i_{1}}$ and $\bm{\beta}^{i_{2}}$ as the size of the intersection between $\mathcal{F}_{k,j}^{i_{1}}$ and $\mathcal{F}_{k,j}^{i_{2}}$ divided by the maximum potential overlap, i.e.,

### v2

@@ 4.3 Vehicle Routing Problem
An instance of CVRP is given by a complete graph $G=(V,A)$, where $V$ consists of a depot node $v_{0}$ and $n$ client nodes each with a corresponding demand $d_{j}$. Moreover, each arc $(v_{j},v_{k})$ has associated costs $c_{jk}$. Lastly, there are $m$ vehicles, each with a capacity of $M$. The goal of the problem is to find at most $m$ routes of minimum costs such that each route starts and ends at the depot, each client is visited exactly once, and the total demand on each route does not exceed the vehicle capacity. To formulate $R_{C}$, we use the Miller-Tucker-Zemlin formulation for the CVRP, which can be found in Equation 13 in the appendix. The decision variables are then denoted by $x_{jk}$ and equal $1$ if arc $(v_{j},v_{k})$ is used in the solution, and 00, otherwise. For this experiment, we consider the parameter vector consisting of the demands $\bm{d}$ and the costs of the arcs from the clients to the depot $\bm{c_{0}}$, i.e., $\bm{\theta}=(\bm{d},\bm{c_{0}})$. The present problem has symmetric costs and consists of 16 clients and 4 vehicles. As a solver for this NP-hard problem, we let the Google OR-Tools heuristic search for a solution for 5 seconds [13]. We aim for explanations for the objective value, and which clients are visited before returning to the depot, i.e., the decision variables $x_{j0}$ for $j\in[n]_{1}$.
As shown in Table 4, the explanation found by CLEMO is significantly more coherent without a considerable loss in accuracy compared to the LR benchmark. We visualize the explanations found by CLEMO in Figure 3, where we first see the solution to the present problem found by Google OR-Tools in gray. Next, the feature contribution of the demands and costs are visualized via node and edge colors respectively. Combined with an overview of the top 10 most contributing features, this shows which features are the key components influencing the objective value as found by the solver. Thus, Figure 3 tells us that the objective value is mainly affected by the distance towards the nodes far from the depot. In Figure 9 in the appendix, we present an additional explanation for the decision variable $x_{20}$.
@@
Feasible region
@@
5 Conclusion & Limitations
In this paper, we propose CLEMO, a sampling-based method that can be used to explain arbitrary exact or heuristic solution algorithms for optimization problems. Our method provides local explanations for the objective value and decision variables of mathematical optimization models. Contrary to existing methods, CLEMO enforces explanations that are coherent with the underlying model structure which enhances transparent decision-making. By applying CLEMO to various optimization problems we have shown that we can find explanations that are significantly more coherent than benchmark explanations generated using LIME without substantially compromising prediction accuracy. At the same time, including coherence losses to CLEMO leads to longer runtimes.
This work focuses on explaining the objective value and decision variables. However, one could easily extend the concept to explanations of other components such as constraint slacks, runtime, optimality gap, etc. For now, CLEMO uses parametric regression models for explanations. Another extension of our work could be to consider other types of interpretable functions such as decision trees. Lastly, CLEMO could be a useful method to explain synergies between ML and optimization models, e.g., in predict-then-optimize models.
@@ A.1 Mathematical Models of The Optimization Problems
In this section, we present the formulation of the Shortest Path problem Equation 12 and the Capacitated Vehicle Routing problem Equation 13. For the latter, we use the Miller-Tucker-Zemlin formulation as described in [19].
@@ A.3 Coherence for Linear Optimization Problems
$\min_{\bm{\beta}_{x_{j}}}\ \sum_{i=0}^{N}w^{i}\|\bm{x}^{i}_{j}-\bm{\beta}_{x_{j}}^{\intercal}\bm{\theta}^{i}\|^{2}\quad j=1,\ldots,p$
fulfill the coherence condition in (3).
@@
$\bm{\beta}_{f}^{*}=(\bm{\Theta}^{\intercal}\bm{W}\bm{\Theta})^{-1}\bm{\Theta}^{\intercal}\bm{W}\bm{y^{f}},$
@@
$=\sum_{j=1}^{p}\hat{c}_{j}\bm{\theta}^{\intercal}\bm{\beta}_{x_{j}},$
@@ A.4.2 Knapsack Problem
Next to the standard deviation of feature contribution, we consider an additional measure for stability, the feature stability index (FSI). This is an adaptation of the variables stability index (VSI) as presented in [35]. The higher this measure, the more the non-zero features found by the different models due to resampling overlap. For a consistent explanation, the overlap should be large. As we apply CLEMO on 10 different datasets for each instance of each type of knapsack problem, we obtain 10 surrogate models given by $\bm{\beta}^{1}_{CL},\dots,\bm{\beta}^{10}_{CL}$. We denote $\mathcal{F}_{k,j}^{i}$ for the set of the top-$k$ most contributing, non-zero features of the $j$-th component of $\bm{\beta}^{i}_{CL}$. We define the $(k,j)$-concordance of two models $\bm{\beta}^{i_{1}}$ and $\bm{\beta}^{i_{2}}$ as the size of the intersection between $\mathcal{F}_{k,j}^{i_{1}}$ and $\mathcal{F}_{k,j}^{i_{2}}$ divided by the maximum potential overlap, i.e.,